60 miles per hour into meters per second
1
Expert's answer
2012-08-30T09:30:20-0400
We know that 1 mile = 1609.344 meters and that 1 hour has 3600 seconds. So,

60 miles per hour
= 60 miles per 3600 seconds
= 60/3600 miles per second
= 1/60 miles per second
= 1/60 * 1609.344 meters per second
= 26,8224 m/s.